23. IVIG infusion: the ordered dose is 2 gm/kg/dose. (IVIG concentration is 2 gm/10 ml 0.9 NS). The nurse runs the IVIG over 6 hours, Patient wt 18 kg

a.) Calculate the correct dose for this patient: ____________

b.) The first ½ hour: run infusion at ¼ rate for 30 minutes – what is that rate? ___________/hr

c.) The second ½ hour: run infusion at ½ rate for 30 minutes – what is that rate? ___________/hr

d.) If the infusion is tolerated increase to the rate of 2 ml/kg/hour until complete. What is that rate: ______ / hour
Answer #1

23) ordered dose of IVIG is 2gm /kg/dose

Available IVIG is 2gm /10ml0.9%NS

Total time of infusion is 6hours

Weight of the patient is 18kg

A) dose for this patient is 2gm/kg/dose

=2×18 /dose

36gm of IVIG per dose

B) infusion rate is 1/4 for 30mins.

Rate = total volume /total time

= 10 ml/ 6hours = 1.66ml/hr

1/4 of 1.66= 0.415 ml /he in 30mins

C) infusion rate is 1/2 for 30mins.

The rate will be = 1/2 ×1.66 = 0.83 ml/hr

D) Rate of infusion is 2ml/ kg/ hour ,if infusion is tolerated.

= 2ml×18= 36ml /hour.